# Iterative Refinement of Project-Level Code Context for Precise Code Generation with Compiler Feedback

### v1

@@ 1 Introduction
Practical code generation for software projects requires using project-specific contexts like classes, methods, and data structures due to modularity, maintenance, and code independence requirements.
@@
However, the repository’s size often exceeds the capacity of the prompt, challenging LLMs to comprehend extensive contexts Liao et al. (2023); Li et al. (2024). Also, studies have demonstrated that irrelevant context in prompts potentially compromises the model performance Shi et al. (2023). Inputting only the task requirements can lead LLMs to overlook project-specific APIs, classes, data structures, or type information specific to a software project’s repository, risking omitting essential logic during code generation Gifany et al. (2013).
As a motivation example, consider the the case shown in Figure 1. In this case, we use the OpenAI GPT-3.5-Turbo API to generate an image-enhancing function by first reducing the notice and then adjusting brightness and contrast. Given the prompt in Figure 1a, the LLM will produce a code snippet of calling a reduce_noise method as depicted in Figure 1b. Although the generated code follows a standard workflow specified in the task prompt, it leads to a compilation error in our application context because reduce_noise is not implemented. This is unsurprising as the prompt does not provide enough project-level context for the LLM, and can be fixed by providing the project-specific function denoising.denoise as context, as illustrated in Figure 1c. While a carefully engineered prompt may resolve this issue, it is not always possible for the user to generate such prompts, and the project context may be too large to fit into the prompt. As we will show later in the paper, such errors are commonly found when applying LLMs to repository-level code generation Li et al. (2024).
@@ 2.2 Error Analysis in Code Generation
As the code is generated, we compile the code and collect the errors reported by the compiler or in testing. Each generated code snippet contains precisely one type of error. If multiple errors are reported by the compiler, the most common error type is selected for analysis. The error distribution revealed that four specific types of errors constitute the majority of all errors encountered. We categorize these errors into: 1) UNDEF, involving Use of Undefined Symbol, 2) API, involving Incorrect Use of APIs, 3) OBJECT, involving Improper Use of an Object, 4) FUNC, involving Runtime or Functional Errors, and 5) OTHER, involving Other Syntax and Semantic Errors. Table 1 presents several errors encountered in compiling and testing. One example is the “UNDEF” error, where a variable AsyncBolt5x0 is referenced but does not exist in the specified module.
@@ 3.3 Retrieval-Augmented Code Generation
Based on the compiler feedback, we aim to retrieve the relevant project-level context from all extracted ones. We implement this by transforming the textual compiler feedback into an SQL query using the ChatGPT. The prompt used is presented in Figure 4(a). Here, several examples of paired compiler feedback and SQL queries are provided as demonstrations for in-context learning [^1] Only one demonstration example is illustrated in Figure 3(b), the full list of examples can be found in Appendix C..
@@
Using this SQL query, the code snippets that involve the implementations of AsyncBolt3 will be returned from our constructed database.
@@
In addition to returning the project-level level retrieved by the SQL query, we also enhance the acquired context by retrieving similar code snippets using dense passage retrieval.
Specifically, given a natural language query $q$, we first convert it into an embedding vector by utilizing an encoder network, as follows:
@@
After generating the query vector, we calculate the cosine similarity between the query and embedding vectors of each context entry $\bm{h}_{c}$. This similarity measure is defined as:
@@ 3.4 Refinement with Compiler Feedback
Figure 3(b) showcases the iterative refinement pipeline. Given the task requirement and partial function $f_{1}$, a semantic retrieval is activated to identify similar functions. Specifically, the function $f_{2}$, which provides equivalent functionality in synchronous scenario, is identified. Utilizing both the retrieved context and the prompt illustrated in Figure 4(b), the language model generates a solution; However, the generated output mistakenly invokes SyncBolt3 due to its intended use in asynchronous scenarios, not aligning with the synchronous scenario in $f_{2}$ The compiler’s feedback highlights this error. With this feedback, ProCoder conducts a hybrid structural and semantic search, leading to the discovery of the correct function $f_{3}$ for asynchronous scenarios. Incorporating the error details and context into the next iteration ensures accurate function invocation. This process goes iteratively until no error is reported by the compiler, resulting in an error-free solution that aligns with the project’s environment.
@@ 5.2 Usefulness of Compiler Feedback
Here, we examine the usefulness of compiler feedback by integrating it into three baseline models: Vanilla, ReACC, and RepoCoder, each considered separately. We conducted experiments using the GPT-3.5-Turbo and subsequently reported the average score across the entire CoderEval dataset, as shown in Table 3. The table provides clear evidence that incorporating compiler feedback yields a notable enhancement in model performance for code generation. Specifically, a comparison between RepoCoder with and without compiler feedback reveals a substantial increase in Pass@$1$ from 36.82 to 38.00. Note that, the RepoCoder+CF here is equivalent to our proposed ProCoder.
@@ 5.4 Error Analysis and Case Study
To thoroughly evaluate ProCoder’s effectiveness, we focused on scenarios where compilation was successful, but execution failed. In the case illustrated in Figure 7, ProCoder incorrectly excludes the mileseconds field and erroneously adds a month field. This error stems from ambiguously stated task requirements and the model’s lack of familiarity with the Structure class’s format, despite its definition being available, resulting in misinterpretation of the intended functionality. Further examples are detailed in Appendix A. This observation inspires us to integrate a comprehensive reference comprising both documentation and web search results to provide explicit usage guidelines in our future work.
@@ 6 Related Work
Automated code generation has a history spanning several decades, with initial endeavors utilizing rule-based systems Backus et al. (1957); Woods (1973) and structured prediction Zelle and Mooney (1996); Zettlemoyer and Collins (2005). In recent years, the development of large-scale language models has led to the emergence of many prominent models in coding tasks. These include open-access models such as DeepSeek Coder Bi et al. (2024), Code Llama Rozière et al. (2023), CodeGen Nijkamp et al. (2022), StarCoder Li et al. (2023), and CodeGeeX2 Zheng et al. (2023), alongside commercial offerings like GPT-3.5 OpenAI (2023a), Gemini Google (2024), and LLM-powered coding assistant tools such as GitHub Copilot Microsoft (2024) and Tabnine Tabnine (2024). These models and tools have demonstrated significant promise in enhancing code generation capabilities.
Context-aware Code Generation.
In the context of project-level code generation, harnessing information beyond the current file has been found useful. The project can be extracted structurally into various forms like relational database Pashakhanloo et al. (2022), API dependency graph Lyu et al. (2021), and class hierarchies Zhang et al. (2021). Shrivastava et al. (2023) proposes a code completion framework using repository-level prompt proposals, and then applying a classifier to classify useful proposals. Lu et al. (2022); Zhang et al. (2023a) propose to use single or multiple levels of retrieval-augmented generation mechanisms for code generation. Yu et al. (2023); Liu et al. (2023); Zhou et al. (2022) proposes benchmarks and datasets for repository-level code synthesis tasks. In our work, we select two code models, namely GPT-3.5 and Code Llama, and retrieve project context for fixing context-related compilation errors, to enhance the precision of code generation.
@@ Appendix A More details about the Error Cases
In this section, we delve into additional cases where ProCoder struggles to generate viable solutions, elucidating the underlying causes to foster further research. We identify two primary factors leading to execution errors despite successful compilation: degenerate solutions, such as overly simplistic or redundant code. This phenomenon is discovered and explored in Zhang et al. (2023b). Figure A1 exemplifies this issue, showcasing a solution that omits necessary validity checks and error handling, neither of which are specified in the task requirements, nor produced by ProCoder’s outputs. Our analysis reveals that degenerate solutions frequently pass compilation, rendering compiler-based verification ineffective.
@@ Appendix B Errors Reported by the Compiler
Appendix C Demonstration Examples of Structural Queries
@@
Appendix D Algorithm for Project Database Construction

### v2

@@ 1 Introduction
Practical code generation for software projects requires using project-specific contexts like classes, methods, and data structures due to modularity, maintenance, and code independence requirements. However, the repository’s size often exceeds the capacity of the prompt, challenging LLMs to comprehend extensive contexts Liao et al. (2023); Li et al. (2024). Also, studies have demonstrated that irrelevant context in prompts potentially compromises the model performance Shi et al. (2023). Inputting only the task requirements can lead LLMs to overlook project-specific APIs, classes, data structures, or type information specific to a software project’s repository, risking omitting essential logic during code generation Gifany et al. (2013).
@@
As a motivation example, consider the case shown in Figure 1. In this case, we use the OpenAI GPT-3.5-Turbo API to generate an image-enhancing function by first reducing the notice and then adjusting brightness and contrast. Given the prompt in Figure 1a, the LLM will produce a code snippet of calling a reduce_noise method as depicted in Figure 1b. Although the generated code follows a standard workflow specified in the task prompt, it leads to a compilation error in our application context because reduce_noise is not implemented. This is unsurprising as the prompt does not provide enough project-level context for the LLM, and can be fixed by providing the project-specific function denoising.denoise as context, as illustrated in Figure 1c. While a carefully engineered prompt may resolve this issue, it is not always possible for the user to generate such prompts, and the project context may be too large to fit into the prompt. As we will show later in the paper, such errors are commonly found when applying LLMs to repository-level code generation Li et al. (2024).
@@ 2.2 Error Analysis in Code Generation
As the code is generated, we compile the code and collect the errors reported by the compiler or in testing. We have generated 10 candidate solutions for a task, comprising 850 solutions for function-level tasks and 1450 for project-level tasks. We report the Pass@10 rate as 53.57% for function-level tasks and 39.73% for project-level tasks. The code that did not pass the test has been taken into error analysis. Each generated code snippet contains precisely one type of error. If multiple errors are reported by the compiler, the most common error type is selected for analysis. The error distribution revealed that four specific types of errors constitute the majority of all errors encountered. We categorize these errors into: 1) UNDEF, involving Use of Undefined Symbol, 2) API, involving Incorrect Use of APIs, 3) OBJECT, involving Improper Use of an Object, 4) FUNC, involving Runtime or Functional Errors, and 5) OTHER, involving Other Syntax and Semantic Errors. Table 1 presents several errors encountered in compiling and testing. One example is the “UNDEF” error, where a variable AsyncBolt5x0 is referenced but does not exist in the specified module.
@@ 3.3 Retrieval-Augmented Code Generation
Based on the compiler feedback, we aim to retrieve the relevant project-level context from all extracted ones. We implement this by transforming the textual compiler feedback into an SQL query using the ChatGPT. The prompt used is presented in Figure 4(a). Here, several examples of paired compiler feedback and SQL queries are provided as demonstrations for in-context learning [^1] Only one demonstration example is illustrated in Figure 3(b), the full list of examples can be found in Appendix C.2..
@@
Using this SQL query, the code snippets that involve the implementations of AsyncBolt3 will be returned from our constructed database. The detailed process of constructing and querying such SQL database is refer to Appendix C.
@@
In addition to returning the project-level level retrieved by the SQL query, we also enhance the acquired context by retrieving similar code snippets using dense passage retrieval. In the initial search round, no compilation error is reported, and ProCoder utilizes the task description string for retrieval. In subsequent searches, ProCoder employs the error report and the corresponding error line.
Given a natural language query $q$, ProCoder first convert it into an embedding vector by utilizing an encoder network, as follows:
@@
After generating the query vector, ProCoder calculate the cosine similarity between the query and embedding vectors of each context entry $\bm{h}_{c}$. This similarity measure is defined as:
@@ 3.4 Refinement with Compiler Feedback
Figure 3(b) showcases the iterative refinement pipeline. Given the task requirement and partial function $f_{1}$, a semantic retrieval is activated to identify similar functions. Specifically, the function $f_{2}$, which provides equivalent functionality in synchronous scenario, is identified. Utilizing both the retrieved context and the prompt illustrated in Figure 4(b), the language model generates a solution; However, the generated output mistakenly invokes SyncBolt3 due to its intended use in asynchronous scenarios, not aligning with the synchronous scenario in $f_{2}$ The compiler’s feedback highlights this error. With this feedback, ProCoder conducts the structural and semantic search, leading to the discovery of the correct function $f_{3}$ for asynchronous scenarios. Incorporating the error details and context into the next iteration ensures accurate function invocation. This process goes iterative until no error is reported by the compiler, resulting in an error-free solution that aligns with the project’s environment. It is noteworthy that ProCoder does not take into account FUNC errors, which arise during execution despite successful compilation. ProCoder focusing on addressing compilation errors, which constitute 66% of the total errors in the context of project-level code generation, as shown in Figure 6.
@@ 4.3 Implementation Details
We have selected GPT-3.5-turbo and Code-Llama 13B for investigation. For GPT-3.5-turbo, the invocation is performed through an online API. For Code-Llama 13B, we utilize its base variant, specifically designed for general code synthesis and understanding. [^2] The model can be accessed via https://huggingface.co/codellama/CodeLlama-13b-hf.
@@ 5.2 Usefulness of Compiler Feedback
Here, we examine the usefulness of compiler feedback by integrating it into three baseline models: Vanilla, ReACC, and RepoCoder, each considered separately. We conducted experiments using the GPT-3.5-Turbo and subsequently reported the average score across the entire CoderEval dataset, as shown in Table 3. The table provides clear evidence that incorporating compiler feedback yields a notable enhancement in model performance for code generation. Specifically, a comparison between RepoCoder with and without compiler feedback reveals a substantial increase in Pass@$1$ from 36.82 to 38.00. More analytic experiments including the ablation study, ProCoder’s performance on function-level code generation, efficiency of static analysis and SQL queries, could be found in Appendix D
@@ 5.4 Error Analysis and Case Study
To thoroughly evaluate ProCoder’s effectiveness, we focused on scenarios where compilation was successful, but execution failed. In the case illustrated in Figure 7, ProCoder incorrectly excludes the microseconds field and erroneously adds a month field. This error stems from ambiguously stated task requirements and the model’s lack of familiarity with the Structure class’s format, despite its definition being available, resulting in misinterpretation of the intended functionality. Further examples are detailed in Appendix A. This observation inspires us to integrate a comprehensive reference comprising both documentation and web search results to provide explicit usage guidelines in our future work.
@@ 6 Related Work
Automated code generation has a history spanning several decades, with initial endeavors utilizing rule-based systems Backus et al. (1957); Woods (1973) and structured prediction Zelle and Mooney (1996); Zettlemoyer and Collins (2005). In recent years, the development of large-scale language models has led to the emergence of many prominent models in coding tasks. These include open-access models such as DeepSeek Coder Bi et al. (2024), Code Llama Rozière et al. (2023), and StarCoder Li et al. (2023), alongside commercial offerings like GPT-3.5 OpenAI (2023a), and GitHub Copilot Microsoft (2024) . These models and tools have demonstrated significant promise in enhancing code generation capabilities.
Project-level Code Generation.
Generating accurate code within a project poses challenges due to the modular design of software engineering, which results in cross-file dependency patterns Parnas (1972). Early work augmented N-gram, RNN, and LSTM models with an additional cache model to track local changes Tu et al. (2014); Hellendoorn and Devanbu (2017). Zhang et al. (2021) incorporate context from parent class to generate comments for the child class. Pashakhanloo et al. (2022) transforms projects into a relational database and proposed a graph walking method to traverse relation path and incorporate all relevant attributes along the way. Lyu et al. (2021) constructs an API dependency graph and utilizes a sequence-to-sequence model that incorporates graph embedding to predict the global-context-aware target code. Zan et al. (2022) first use private API documentation to improve code generation, and Zan et al. (2023) investigates how various components of API documentation influence the prediction accuracy. Shrivastava et al. (2023) proposes a code completion framework using repository-level prompt proposals, and then applying a classifier to classify useful proposals. Lu et al. (2022); Zhang et al. (2023a) propose to use single or multiple levels of retrieval-augmented generation mechanisms for code generation. Liao et al. (2023) proposes A3-CodGen, which utilize local, global, and third-party-library information for better context retrieval. Yu et al. (2023); Liu et al. (2023); Zhou et al. (2022); Ding et al. (2024) proposes benchmarks and datasets for repository-level code synthesis tasks.
Post-processing of LLMs for Code Generation
To identify correct code generated by LLM, researchers employ post-processing techniques to further rank and filter the generated code.  Inala et al. (2022) trained a fault-aware neural ranker that ranks multiple code samples based on compilation feedback. AlphaCode Li et al. (2022) and Shi et al. Shi et al. (2022) employ clustering and filtering methods based on the execution output of the generated program. Tsai et al. (2023) proposes to retrieve compilation error patches from a database of human expert guidance documents for RTL assembly.  Zhang et al. (2023c) rerank LLM outputs based on back translation. SelfEdit Zhang et al. (2023b) employs a generate-and-edit approach that utilizes execution results to improve the competitive programming task code quality.  Chen et al. (2023) propose a rubber duck debugging mechanism without human feedback, which identifies mistakes by investigating the execution results and explaining the generated code in natural language.
@@ Appendix A More details about the Error Cases
In this section, we delve into additional cases where ProCoder struggles to generate viable solutions, elucidating the underlying causes to foster further research. We identify two primary factors leading to execution errors despite successful compilation: degenerate solutions, such as overly simplistic or redundant code. This phenomenon is discovered and explored in Zhang et al. (2023c). Figure A1 exemplifies this issue, showcasing a solution that omits necessary validity checks and error handling, neither of which are specified in the task requirements, nor produced by ProCoder’s outputs. Our analysis reveals that degenerate solutions frequently pass compilation, rendering compiler-based verification ineffective.
@@ Appendix B Errors Reported by the Compiler
Table A2: Tables pre-computed by ProCoder for error correction.
Table Name Description Element Example
M Stores the a module and its hierarchy in project. tests.unit.async_.work.__init__
M_C Stores a module and a class inside the module Module neo4j._codec.packstream.v1.__init__, Class PackableBuffer
M_C_CF Stores a class, its parent module, and its member functions. Module neo4j.time.__init__, Class Clock, Function local_offset
M_C_V Stores a class variable, its parent class and module. Module neo4j._sync.io.tmphhoug1of, Class Bolt, Variable is_reset
M_GF Store a global function and its parent module. Module neo4j.time._arithmetic, Function nano_add
M_GV Stores a global variable and its parent module. neo4j.__init__, Global Variable TRUST_SYSTEM_CA_SIGNED_CERTIFICATES
@@
Appendix C Technical Details of Structural Query System
C.1 The Design Principle
In ProCoder, we employ the Structural Query Language (SQL) to perform structured queries on code repositories. The SQL syntax is derived from CodeQL, which is an analysis tool designed for vulnerability mining through structured queries, capable of conducting complex analyses on code repositories. We extracted 11 error codes from compiler (shown in Table A1), each corresponding to one of four error categories as in Section 2.2. Two authors manually inspected these error codes, and constructed 6 SQL data tables from the project graph, based on the characteristics of the errors. All tables are presented in Table  A2. Whenever an error is reported by the compiler, the LLM generates an SQL query statement for it, where the modules and functions to be queried are extracted using a tree-sitter-based parser  [^1] https://github.com/tree-sitter/tree-sitter-ql. In the ProCoder with iteration rounds set to 3, 93.2% of the SQL queries generated by the LLM pass the parsing utility, and these queries are similar to the demonstration example SQL queries. Subsequently, according to the entity from the query, it is mapped to one of the six SQL data tables. We store these six tables in a relational database for flexible retrieval.
C.2 Demonstration Examples of Structural Queries
@@
Appendix D More Analytic Experiments
D.1 Ablation on Components
Table A4: Pass rates of ProCoder on Class Runnable Test.
Data Split Class Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 28.00 44.92 49.09
ProCoder w/o CF and SQL (RepoCoder) 35.45 40.46 41.82
w/ CF 30.00 42.58 45.45
w/ CF+SQL 30.36 44.61 49.09
w/ CF+Semantic 31.45 45.00 47.27
Table A5: Pass rates of ProCoder on File Runnable Test.
Data Split File Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 30.29 43.58 47.06
ProCoder w/o CF and SQL (RepoCoder) 29.41 34.61 36.76
w/ CF 26.03 37.41 42.65
w/ CF+SQL 26.76 37.68 39.71
w/ CF+Semantic 27.35 39.90 45.45
Table A6: Pass rates of ProCoder on Project Runnable Test.
Data Split Project Runnable
Method Pass@1 Pass@5 Pass@10
ProCoder 21.30 36.73 39.13
ProCoder w/o CF and SQL (RepoCoder) 16.96 19.57 21.74
w/ CF 14.35 25.62 30.43
w/ CF+SQL 15.65 25.12 30.43
w/ CF+Semantic 17.83 28.50 34.78
To explore the impact of each component within ProCoder on overall performance, we remove various components from ProCoder, including the compiler, SQL retriever, and semantic retriever. Notably, the configuration that excludes the compiler and SQL, relying solely on semantic retrieval, is known as RepoCoder Zhang et al. (2023a).
Tables A4, A5, and A6 demonstrate the Pass Rates of ProCoder across CoderEval-python. The data indicates that incorporating Compiler Feedback does improve accuracy, although not markedly substantial. The reason is that highlighting compilation errors alone does not provide related context for resolving them, which is important in project-level coding problems. Also, compared to RepoCoder—which relies solely on Semantic Retrieval, integrating Compiler Feedback with contextual information results in a dramatic performance improvement.
D.2 Varying the Number of Iterations
Table A7: Pass rates of ProCoder on varying iterations.
Iteration Class Level File Level Project Level
$i=0$ 34.55 30.88 13.04
$i=3$ 49.09 47.06 39.13
$i=10$ 47.27 45.59 39.13
To investigate whether iteration could continuously improve performance, we increased the number of iterations to ten. As shown in Table A7, with more iterations, the performance of ProCoder reaches a plateau. This indicates that there are still errors that cannot got repaired by ProCoder, leaving for further research.
D.3 ProCoder on Function-level Code Generation
Table A8: The performance of CoCoGen across varying iterations on the HumanEval benchmark. (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 61.52 80.56 83.15 40 2.38%
CF $i=1$ 71.46 82.46 85.2 7 0.42%
CF $i=2$ 71.4 82.56 85.98 3 0.18%
CF $i=3$ 71.65 82.49 85.36 3 0.18%
Table A9: The performance of CoCoGen across varying iterations on the MBPP benchmark. (CE: Compilation Errors, CE%: their proportion among all generated solutions)
Pass@1 Pass@5 Pass@10 CE CE(%)
Direct 49.71 58.50 60.81 185 1.90%
CF $i=1$ 52.95 59.58 61.85 11 0.11%
CF $i=2$ 53.06 59.87 62.01 9 0.09%
CF $i=3$ 52.79 59.80 62.09 7 0.07%
To assess the efficacy of ProCoder in function-level code generation, we evaluated it on two commonly used datasets: HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). We retained the compiler feedback model in ProCoder, and conduct tests in a 0-shot scenario, generating 10 code samples for each of the 163 test cases for HumanEval and 974 cases for MBPP, totaling 11,370 samples. The generation and refinement loop iterates 3 times. As shown in Table A8 and Table A9.
The data presented in the table illustrates that compilation feedback improves the accuracy of the generated code, although its efficacy is somewhat limited. We manually review 7 compilation errors not resolved in HumanEval after the first iteration, and find that 6 out of 7 instances were due to LLM erroneously generating code segments enclosed in Markdown code block markers (“‘), resulting in compilation failures. Additionally, another instance of compilation error was identified in the following generated solution:
def is_simple_power(x, n): return x > 0 and (x == 1 or (n != 1 and x == n**int(round(math.log(x, n))))
This code snippet contains five left parentheses and four right parentheses in the return statement, causing the syntax error. Interestingly, despite the compiler has indicated this syntax error, it was not rectified over three iterations.
In MBPP, we manually observe the 7 errors not fixed after the last iteration, and identifies that two test tasks were responsible for all 7 errors: one named "sum" contributed to six errors, and another called "month_season" resulted in one error. We detail these error cases as follows:
def month_season(month, day): seasons = "spring": [(3, 20), (6, 20)], "summer": [(6, 21), (9, 22)], "autumn": [(9, 23), (12, 20)], "winter": [(12, 21), (3, 19)] for season, (start, end) in seasons.items(): if (month == start[0] and day >= start[1]) or (month == end[0] and day <= end[1]): return season return "Invalid input"
The error arises because ’start’ and ’end’ are not included in seasons.items, which appears strange. We speculate that this issue stems from the presence of functions with the same name within the training set.
def sum(a, b): common_divisors = [i for i in range(1, min(a, b) + 1) if a % i == 0 and b % i == 0] return sum(common_divisors) if common_divisors else 0
The root cause of this error lies in the model’s confusion when a user-defined function name coincides with that of a system library function sum, representing a category of potential issues.
D.4 Efficiency of Static Analysis
Utilizing static analysis tools to inspect code typically results in increased latency. To investigate whether the latency impacts the usability of ProCoder, we measure its latency at each test case. We have observed that in experiments, the speed of static analysis and structural query is relatively fast. It is because the semantic checker (i.e., pylint), only analyzes the current file and its dependent files. The average latency was recorded at 1.27 seconds, with a minimum of 0.359 seconds and a maximum of 6.984 seconds. We deem this latency to be acceptable.
Appendix E Algorithm for Project Database Construction